Calculate the change in internal energy of 3.00 mol of helium gas when its temperature is
increased by 2.00 K.
1
Expert's answer
2021-08-09T11:22:34-0400

Internal energy

"U=\frac{3}{2}nRT"

"∆U=\frac{3}{2}nR∆T"

Put value


"∆U=\frac{3}{2}\times3\times8.31\times2=74.79J"